A student has computed that it takes an average (mean) of 16 minutes with a standard deviation of 4 minutes to drive from home, park the car, and walk to an early morning class. Another day it only took the student 12 minutes to get to class. What is the measurement in standard units
Answer by stanbon(75887)

A student has computed that it takes an average (mean) of 16 minutes with a standard deviation of 4 minutes to drive from home, park the car, and walk to an early morning class. Another day it only took the student 12 minutes to get to class. What is the measurement in standard units
----
z(12) = (12-16)/4 = -4/4 = -1
=================================
Ans: One standard deviation below the mean.
=================================
Cheers,
Stan H.